OpenStudy (anonymous):

solve the equation for principal values of X. Express solution in degrees cos x= 3cos x-2

OpenStudy (anonymous):

\[2\cos(x)-2=0\] is a good start

OpenStudy (fibonaccichick666):

can you put some parentheses in so the eq is clear please

OpenStudy (anonymous):

that was whats on the worksheet hun

OpenStudy (anonymous):

so next it would be 2 cos x= 2?

OpenStudy (anonymous):

yes

OpenStudy (fibonaccichick666):

I'm sure there are some parentheses around the inside of the cosine, ie is it cosine(x-2) or is it cos(x)-2

OpenStudy (anonymous):

and then divided by two?

OpenStudy (anonymous):

so cos x=1?

OpenStudy (anonymous):

so im confused dose x= 0, 180, 360, or 540 degrees?

OpenStudy (freckles):

cos(180)=-1 not 1 cos(540)=-1 not 1

OpenStudy (anonymous):

so its 360 or 0?

OpenStudy (freckles):

well there are infinitely many solutions

OpenStudy (freckles):

also what does principal values mean here?

OpenStudy (anonymous):

yea i got that but my teacher wants one

OpenStudy (anonymous):

i have no idea, that's one of the reasons Im confused

OpenStudy (freckles):

can't help you if you don't tell me what principal values mean

OpenStudy (freckles):

it is probably 0 though

OpenStudy (anonymous):

a principal value is a value along one branch of the function. but ill go with 0

OpenStudy (freckles):

@misty1212 what the heck does that definition mean? what is a branch of a function?

OpenStudy (anonymous):

haha 0 was correct i checked with my older brother no worries

OpenStudy (freckles):

I still want to know what branch of a function means

OpenStudy (freckles):

lol

OpenStudy (anonymous):

its a branch of a function. if you graph a function a piece between two points on the graph would be a branch of the function, at least that's how i would look at it logically then again math is rarely logical

OpenStudy (freckles):

Well it is definitely suppose to be logical. But I think you meant this for the definition of principal values: For sin(x)=a, it is a number x between -90 deg and 90 deg (including endpoints). For cos(x)=a, it is a number x between 0 and 180 deg (including endpoints).

OpenStudy (misty1212):

probably means the answer on the interval \[[0,2\pi)\]

OpenStudy (misty1212):

so in this case you pick \(x=0\)